# Benchmarking Federated Learning for Semantic Datasets: Federated Scene Graph Generation

### v1

@@ Abstract
Federated learning (FL) has recently garnered attention as a data-decentralized training framework that enables the learning of deep models from locally distributed samples while keeping data privacy. Built upon the framework, immense efforts have been made to establish FL benchmarks, which provide rigorous evaluation settings that control data heterogeneity across clients. Prior efforts have mainly focused on handling relatively simple classification tasks, where each sample is annotated with a one-hot label, such as MNIST, CIFAR, LEAF benchmark, etc. However, little attention has been paid to demonstrating an FL benchmark that handles complicated semantics, where each sample encompasses diverse semantic information from multiple labels, such as Panoptic Scene Graph Generation (PSG) with objects, subjects, and relations between them. Because the existing benchmark is designed to distribute data in a narrow view of a single semantic, e.g., a one-hot label, managing the complicated semantic heterogeneity across clients when formalizing FL benchmarks is non-trivial. In this paper, we propose a benchmark process to establish an FL benchmark with controllable semantic heterogeneity across clients: two key steps are i) data clustering with semantics and ii) data distributing via controllable semantic heterogeneity across clients. As a proof of concept, we first construct a federated PSG benchmark, demonstrating the efficacy of the existing PSG methods in an FL setting with controllable semantic heterogeneity of scene graphs. We also present the effectiveness of our benchmark by applying robust federated learning algorithms to data heterogeneity to show increased performance. Our code is available at https://github.com/Seung-B/FL-PSG.
Index Terms:
Scene Graph Generation, Panoptic Scene Graph Generation, Federated Learning, Distributed Learning, Data Privacy, Benchmark
I Introduction
footnotetext: Equal Contributionfootnotetext: Corresponding Authorfootnotetext: This work has been submitted to the IEEE for possible publication.
@@ I Introduction
Federated learning (FL) has drawn great attention as a key framework for enabling decentralized training of deep models from the distributed private data to numerous edge clients. The FL framework communicates the model parameters between the clients and the server; to keep the distributed local data private, the server cannot access data samples of clients [1]. This property of FL that preserves data privacy makes it more crucial when deep models for tasks with license- or privacy-sensitive data, such as clinical data from medical institutions, private information from electronic edge devices, licensed contents from providers, and broadcasting stations, etc.
Along with the rapid advancement of the algorithmic development of FL, great efforts have been dedicated to constructing FL benchmarks that allow reliable and rigorous evaluations for the existing FL methods. The existing FL benchmarks mostly rely on the existing datasets, including MNIST [2], CIFAR [3], CelebA [4], Twitter [5], etc, then building upon them, researchers focus on devising a decentralized training setting with controllable factors, such as data heterogeneity across clients, number of clients, ratio of participating clients, number of the maximum communication rounds, etc.
Figure 1: Examples of Single-semantic (Left: image classification) and Multi-semantic (Right: scene graph generation) tasks.
@@
Among the factors of FL settings, data heterogeneity works as the most crucial factor that obviously exhibits the efficacy of different FL algorithms; when the data distribution strongly deviates across clients, a federation of local models typically fails with drastic performance drops [6]. To construct a controllable and rigorous benchmark process with data heterogeneity, researchers intentionally diversify the prior distribution of one-hot label of samples across clients via random sampling of the priors from Dirichlet distribution [7, 8], or shard- or chunk-wise assignment of data [1].
@@
Herein, we want to point out two key limitations of the existing FL benchmarks. Firstly, the current benchmarks mostly handle simple classification or regression tasks, where each sample is paired with a one-hot label or a single target value. However, deep training tasks currently become far beyond mere classification or recognition and consider highly complicated jobs to understand in-depth semantic information hidden in the given data sample, e.g., generating realistic samples [9, 10], finding similar samples [11, 12, 13] from the prompts, or answering the queries about actions and objects shown in photos [14, 15]. Extending the current FL benchmark process to tasks that consider complicated semantics is necessary.
@@
Second, there does not exist a task-agnostic FL benchmark process that devises controllable semantic heterogeneity of data across clients. In a simple image classification FL benchmark, for instance, the existing process focuses on a single semantic, i.e., the class label, and deviates the label distribution across clients to devise data heterogeneity (referring to the left image of Fig. 1). In contrast, devising data heterogeneity is non-trivial when considering vision tasks, where each sample contains multiple semantics. As shown in the right image of Fig. 1, Scene Graph Generation, which is a core vision task for understanding the complicated semantics of a given image, a single sample bears multiple objects (‘cat’, ‘keyboard’, ‘dog’ and ‘truck’), predicates (‘looking at’ and ‘sitting on’) and relations (‘cat’ $\rightarrow$ ‘keyboard’, ‘dog’ $\rightarrow$ ‘truck’). Partitioning samples with such complex semantics into clients while controlling the semantic heterogeneity remains unexplored.
@@
In this study, we propose an FL benchmark process that enables the evaluation of FL algorithms on multi-semantic datasets while controlling semantic heterogeneity. To break the aforementioned limitations, our process encompasses two key steps: i) discovering the semantic clusters by utilizing the collection of multiple annotations called ‘category tensor’. ii) distributing data samples to multiple clients by considering the heterogeneity between the different semantic clusters.
As a proof of concept, we aim to construct the FL benchmark for the Panoptic Scene Graph Generation (PSG) task, a family of scene graph generation tasks that utilize the panoptic segmentation, i.e., segmented objects, not the bounding boxes. And classic scene graph generation tasks use a bounding box-based object grounding. However, the bounding box contains more noise than the pixel-level object. In scene graph generation, inferring relationships between objects is crucial. Panoptic segmentation provides precise boundary information, allowing better computation of distances or overlaps between objects. We select the PSG task because discovering scene graphs not only directly links to the fundamental understanding of visual perceptions but also works as a key module in bridging vision and language [16].
To the best of our knowledge, our work first attempts to establish the FL benchmarks for the PSG and provides the evaluation results that demonstrate the effectiveness of the existing PSG baselines in decentralized training settings. The simulation results reveal that the methods tailored to tackle the long-tailed problem in the PSG task, where some objects and predicates are more dominant than others, are robust in handling semantic heterogeneity in FL.
II Related Works
Figure 2: The overview of federated learning process. (a): The server distributes a global model to clients. (b): Each client initializes the local model with the distributed one and trains a neural network using their dataset (c): Clients upload the locally trained model parameters to the server. Then, the server averages the aggregated models.
II-A Federated Learning (FL) and Benchmarks
FL has emerged as a pivotal framework for training deep learning models in a decentralized setting, enabling the preservation of data privacy for clients. Clients maintain their private data and do not transfer sensitive information to a central server. However, data privacy causes heterogeneity problems, where clients have different data distributions, and the server can not control clients. Therefore, the trained model of each client diverges, which is a crucial factor for convergence and performance in FL. Therefore, researchers mainly have been dedicated to handling the case with a substantial heterogeneity of data across clients, resulting in the diverse strategies, including FedAvgM [17], FedProx [18], SCAFFOLD [19], FedOpt [20], FedDyn [21], and FedGF [22].
Researchers have suggested a variety of FL datasets for various tasks to establish thorough evaluation of the FL algorithms ranging from image classification [5, 23], natural language processing (NLP) [5, 24], audio emotion recognition [25], multimodal learning [26], to graph-based learning [27]. For organizing the agreed framework for the FL settings, several FL testing environments have been publicly researched and released, including Flower [28], FedML [29], FedScale [30], and FedLab [31].
@@ II-A Federated Learning (FL) and Benchmarks
Regardless of the datasets and environments, the most crucial factor in FL evaluation is demonstrating the effectiveness of methods with strong data heterogeneity across clients. When each sample contains a simple semantic, such as a single target label, unified strategies exist to impose heterogeneity across clients by diversifying the prior distribution of the target label. Specifically, two main strategies include i) sampling the prior distribution of each client from Dirichlet distribution [21], and ii) chucking per-class data samples into multiple shards, where a fixed number of shards are allocated to each client, resulting in heterogeneity between clients [32, 33].
However, when each sample bears multiple semantics, the current methods are not easily extended, so many works rely on random splits into clients, which cannot impose heterogeneity across clients. We suggest a benchmark process with well-controlled semantic heterogeneity across clients for fully evaluating deep models in FL settings with multi-semantics. One recent work has suggested a benchmark called FedNLP [24] to impose semantic heterogeneity across clients, particularly in the NLP field, but it relies on the pretrained language model to discover semantic clusters and cannot be extended to vision tasks. In contrast, our benchmark does not rely on extra pretrained models and considers the first-ever developed FL panoptic scene graph generation (PSG) testing environment.
II-B Panoptic Scene Graph Generation (PSG)
Scene graphs are crucial for scene understanding in computer vision tasks, representing objects (nodes), which are represented by bounding boxes or pixel-wise segmentation, and predicates (relationships, edges) in a graph structure. Predicting the bounding boxes and relationships between bounding boxes constitute scene graph generation. Recently, Conditional Random Field (CRF) [34], TransE [35, 36], CNN [37, 38], RNN/LSTM [39, 40, 41, 42, 43], GCN [44, 45, 46] based scene graph generation methods were studied. Subsequently, the PSG task has been proposed [47], which delves deeper into the scene graph generation by using panoptic segmentation masks instead of bounding boxes. The difference between PSG and classic scene graph generation is that PSG uses panoptic segmentation [48] masks rather than bounding boxes. The difference between the two tasks can be expressed as follows:
$\text{Pr}(G|I)=\text{Pr}(B,S|I),$ (1)
$\text{Pr}(G|I)=\text{Pr}(M,S|I),$ (2)
where $I\in\mathbb{R}^{H\times W\times 3}$ is the input image, and $B$ and $M$ means bounding boxes and segmented masks, respectively. $S$ consists of the labels and their relations. And $G$ is the scene graph. Therefore, existing scene graph generation methods can be applied to PSG tasks. Research on PSG [49, 50, 51, 52, 53] has been diverse and rapidly advancing recently.
Long-tailed Problem: The scene graph generation tasks face the long-tailed problem [54]. Positional relationships among objects constitute the majority of the predicates, leading to a visual relationship complexity of $\mathcal{O}(N^{2}R)$ for $N$ objects and $R$ predicates [16]. This exacerbates the long-tailed problem in SGG datasets, prompting various recent approaches [55, 56, 57, 58, 59, 50, 49, 60, 51] have been proposed to address this issue.
@@ II-B Panoptic Scene Graph Generation (PSG)
To perform challenging vision tasks such as PSG, having more data typically leads to training better models. However, in reality, the photos held by different clients are unlikely to be similar, and collecting such data on the server for training poses a threat to data privacy. Despite this, there have been no attempts to apply FL to the PSG task, highlighting the necessity for this research. Furthermore, the data heterogeneity issue in FL appears similar to the long-tailed problem in the scene graph generation tasks. In addition, the long-tail problem is similarly dealt with in FL [61, 62].
III Background
III-A Federated Learning
We briefly introduce the preliminaries of federated learning (FL) by focusing on the foundational baseline, i.e., FedAvg, [1]. FL settings contain a single server and $K$ clients. The training process consists of iterative rounds, where the server and clients communicate the model parameters to each other. At each round $t$, a server initiates the round by broadcasting a global model $w^{t}$ to all clients. Each client then performs local training with its own data to obtain the locally trained model, i.e., $w_{k}^{t}$, where $k$ is the client index. After that, the server aggregates the locally trained model to compute the average model, which works as the global model of the next round:
$w^{t+1}=\sum^{K}_{k=1}\frac{n_{k}}{n}w^{t}_{k},$ (3)
where $n_{k}$ is the number of data samples on client $k$, and $n$ is the total data samples across all clients. Notably, our formulation assumes full participation of all clients, but it can be extended to partial participation by letting a subset of clients participate in the aggregation for each round.
III-B Formalizing Heterogeneous Settings
Figure 3: Category Tensor K-means Clustering Pipeline. (a) We make a graph using the objects, subjects, and predicates of each image. (b) Then, we map each label into the super-classes of fine-grained labels. (c) Categorised relations (subject, object, and predicate) are converted into the category tensor. (d) We transform every input image into a category tensor and perform K-Means Clustering.
The heterogeneity of distributed data is a critical factor for FL. When data distribution across clients is homogeneously, then the data distribution is identical for all clients, the so-called independent and identically distributed (IID) case. Otherwise, when the distribution diversifies across clients, we call it the non-IID case. We here describe the existing methods for establishing the non-IID cases in FL.
Label-based partition: Label-based data partitioning is the most widely-used approach where the dataset is distributed according to the label of samples. For instance, shard-based partitioning [1] divides the dataset into shards with each shard containing data samples of one or a few classes. Each client receives one or more shards randomly, resulting in each client owning data samples that are biased toward specific classes. Dirichlet distribution-based partitioning [21] offers a controllable and flexible method for simulating non-IID data distributions. The Dirichlet distribution is parameterized by a concentration parameter $\alpha$; as $\alpha$ gets close to 0, the sampled prior distribution is biased toward specific classes, in contrast, as $\alpha$ gets close to $\infty$, the sampled prior distribution tends to be uniform over all classes.
Feature- or attribute-based partition: A feature commonly refers to an attribute of data samples. For example, when predicting an outcome based on the values from specific sensors, the location of the sensor can be its feature. In such cases, non-IID data can be constructed based on the location. The data features among clients may be completely non-overlapped, partial-overlapped, or full-overlapped [63]. Non-overlapped features refer to cases where the features in the data are different from each other. For instance, some data might include ‘gender’ and ‘age’ as features, while other data might include ‘height’ and ‘weight’. Partial-overlapped features refer to cases where the data shares some, but not all, features. For example, when taking photos of an object from the left and right sides, the images capture the same object but are not identical. Fully overlapped features refer to cases when the data features are completely identical, which is the most common scenario. Based on the degree of feature overlap, we can dynamically construct data heterogeneity.
Temporal partition: Temporal partitioning leverages the temporal and spatial variability of data to construct non-IID datasets. For instance, when utilizing stock market data, data heterogeneity can be created by assigning data from specific periods to individual clients.
In brief, feature-based and temporal partitioning are strongly related to the given task or datasets, so the applicability to a wide range of FL scenarios is severely limited. Therefore, most of the works with algorithmic developments of FL rely on the label-based partitioning method.
IV A Benchmark Process for FL with Multi-Semantic Datasets
@@
For a given multi-semantic dataset, each data sample contains multiple annotations, i.e., $(x,\mathcal{Y})\in\mathcal{D}$, where $x$ is an input, $\mathcal{Y}=\{y_{1},\cdots,y_{M}\}$ is a multi-semantic label, $M$ is the possible number of labels for each data sample, and $\mathcal{D}$ represents the dataset. Here, we introduce our benchmark process to distribute the multi-semantic data samples into $K$ multiple clients with controllable semantic heterogeneity. The key steps are twofold: i) discovering data clusters with different semantics and ii) data partitioning with controllable semantic heterogeneity across clients.
IV-A Discovering Data Clusters: $K$-means Clustering of Category Tensor
@@
For a given multi-semantic $\mathcal{Y}$, we transform it to category tensor $\mathcal{F}$ by allocating each label $y_{i}$ into an orthogonal axis of the tensor, i.e., $\mathcal{F}(\mathcal{Y})\in\mathbb{R}^{N_{1}\times\cdots\times N_{L}}$, where there are $N_{1},\cdots,N_{L}$ possible categories for each respective label of $\mathcal{Y}$. We then apply $K$-means Clustering on the collection of $\mathcal{F(Y)}_{1}^{|\mathcal{D}|}$ of overall dataset:
@@
where $n$ is the discovered number of clusters that can be determined depending on the dataset and $\mathcal{C}_{i}$ indicates the collection of samples assigned to $i$-th cluster. With the obtained clustering, we can transform a data $(x,\mathcal{Y})$ into $(x,\mathcal{C})$ to impose the cluster label $\mathcal{C}$ with semantic information, which is a one-hot label with $1$ for the assigned cluster. Through this clustering, we can perform the label-based partition while fully utilizing the multi-semantic information of each data sample with cluster label $\mathcal{C}$. We present the overall pipeline to Fig 3.
IV-B Data Partition with Semantic Heterogeneity
We acquire $n$ clusters from (4). It trivially raises the issue that the clusters are not evenly distributed, so the number of samples assigned to each cluster would deviate for different clusters, i.e., cluster imbalance. The cluster imbalance prevents rigorous evaluations of FL to handle semantic heterogeneity because a model becomes overfitted to dominant clusters without balanced training across different semantics. The cluster imbalance stems from the long-tailed problem, which is a key challenge in scene graph generation datasets. In other words, we have to create data heterogeneity for FL, which makes it difficult to distinguish from the long-tailed problem. If the amount of data in each cluster is equalized, the long-tailed problem may be partially alleviated. Furthermore, considering the federated learning scenario, this cluster imbalance is likely to bias the update of the global model in the update direction of users belonging to the dominant cluster. That is, it causes overfitting to a dominant cluster, which makes it difficult to closely compare each method. Consequently, we need to equalize the data quantity of each cluster: $\hat{\mathcal{C}}_{k}=\textnormal{Sample}(\mathcal{C}_{k},m)$, for all $1\leq k\leq n$, where $m=\min_{k\in[n]}\{|\mathcal{C}_{k}|\}$, $|\mathcal{C}_{k}|$ is the cardinality of the $k$-th cluster $\mathcal{C}_{k}$, and $\textnormal{Sample}(\mathcal{C}_{k},m)$ denotes a function that randomly selects $m$ data samples from cluster $\mathcal{C}_{k}$.
Now, we are ready to apply label-based partition to these clusters, enabling us to impose semantic heterogeneity. Our benchmark suggests two partition strategies as follows.

### v2

@@ Abstract
Federated learning (FL) enables decentralized training while preserving data privacy, yet existing FL benchmarks address relatively simple classification tasks, where each sample is annotated with a one-hot label. However, little attention has been paid to demonstrating an FL benchmark that handles complicated semantics, where each sample encompasses diverse semantic information, such as relations between objects. Because the existing benchmarks are designed to distribute data in a narrow view of a single semantic, managing the complicated semantic heterogeneity across clients when formalizing FL benchmarks is non-trivial. In this paper, we propose a benchmark process to establish an FL benchmark with controllable semantic heterogeneity across clients: two key steps are (i) data clustering with semantics and (ii) data distributing via controllable semantic heterogeneity across clients. As a proof of concept, we construct a federated PSG benchmark, demonstrating the efficacy of the existing PSG methods in an FL setting with controllable semantic heterogeneity of scene graphs. We also present the effectiveness of our benchmark by applying robust federated learning algorithms to data heterogeneity to show increased performance. To our knowledge, this is the first benchmark framework that enables federated learning and its evaluation for multi-semantic vision tasks under the controlled semantic heterogeneity. Our code is available at https://github.com/Seung-B/FL-PSG.
keywords:
[…]
1 Introduction
@@
Federated learning (FL) has drawn considerable attention as a key framework to enable decentralized deep model training from private data of numerous clients. The FL framework communicates the model parameters between the clients and the server; to keep the distributed local data private, the server cannot access data samples of clients [23]. The property that FL preserves data privacy makes it more crucial when deep models handle license- or privacy-sensitive data, e.g., clinical data from medical institutions, licensed content from providers, and broadcasting stations.
@@
Along with the rapid algorithmic development of FL, significant efforts have been dedicated to constructing FL benchmarks that enable reliable and rigorous evaluations of methods. The existing FL benchmarks mostly rely on the existing datasets, such as CIFAR [13] and Twitter [3], etc. Therefore researchers focus on devising a decentralized training setting with controllable factors, such as data heterogeneity across clients, number of clients, and participation ratio. Among the factors of FL settings, data heterogeneity works as the most crucial factor that exhibits the efficacy of different FL algorithms; when the data distribution strongly deviates across clients, a federation of local models typically fail with drastic performance drops [16]. Researchers have mainly focused on data heterogeneity constructed by a single label, which is straightforward. Therefore, prior works diversify the distribution across clients via Dirichlet distribution [1] or shard- or chunk-wise assignment of data [23].
@@
Herein, we point out two key limitations of the existing FL benchmarks. Firstly, the current benchmarks mainly handle classification or regression tasks, where each sample consists of a single label. However, deep learning tasks are becoming far beyond classification or recognition, and complicated jobs are being considered to understand in-depth semantic information. Therefore, extending the current FL benchmark process to complicated semantics is necessary.
Figure 1: Examples of single-semantic (Left: classification) and multi-semantic (Middle: scene graph generation) tasks, and the overview of the federated learning process (Right).
@@
Second, there does not exist a task-agnostic FL benchmark process that devises controllable semantic heterogeneity. The classification task of FL focuses on a single semantic (the class label) in the left of 1 and constructs heterogeneity based on the label. In contrast, devising heterogeneity is non-trivial when each sample contains multiple semantics. As shown in the middle of 1, Scene Graph Generation, which understands the complicated semantics of an image, a single image contains multiple objects (‘cat’, ‘keyboard’, ‘dog’ and ‘truck’), predicates (‘looking at’ and ‘sitting on’) and relations (‘cat’ $\rightarrow$ ‘keyboard’, ‘dog’ $\rightarrow$ ‘truck’). It remains unexplored to construct the heterogeneity with controllable and complex semantics.
@@
In this study, we propose a comprehensive FL benchmark process for evaluating FL algorithms on multi-semantic datasets while controlling semantic heterogeneity. To break the limitations of existing studies, our process encompasses two key steps: (i) discovering the semantic clusters by utilizing the collection of multiple annotations. and (ii) distributing data samples to multiple clients by considering the heterogeneity. The simulation results reveal that the methods tailored to tackle the long-tailed problem in the Panoptic Scene Graph Generation (PSG) task, where some objects and predicates are more dominant than others, are robust in handling semantic heterogeneity in FL.
2 Related Works
FL has emerged as a framework for training deep learning models in a decentralized setting, enabling the preservation of data privacy for clients. We briefly introduce preliminaries by focusing on the foundational baseline, i.e., FedAvg, [23]. FL setting contains the single server and $K$ clients. The training process proceeds iteratively in rounds. Each round follows the three-step procedure shown on the right of 1, summarized as follows: (a): The server distributes a global model to clients. (b): Each client initializes the local model with the distributed one and trains a neural network using their dataset (c): Clients upload the locally trained model parameters to the server. Then, the server averages the aggregated models and performs the step (a) again. The aggregation step is represented as $w^{t+1}=\sum^{S_{t}}_{k=1}\frac{n_{k}}{n}w^{t}_{k}$, where $n_{k}$ is the number of data samples on client $k$, $S_{t}$ is the set of selected clients at round $t$, and $n$ is the total number of data samples across $S_{t}$ clients. When $S_{t}$ equals $K$, all the clients participate in the aggregation step, and in the case of $S_{t}<K$, partial clients participate in the aggregation step.
In FL systems, each user has independent data with different distributions, which can cause the trained local models to diverge, negatively impacting the performance and convergence of FL. Therefore, researchers have mainly been dedicated to handling the case with a substantial heterogeneity of data across clients, resulting in diverse strategies. For example, FedAvgM [8] leverages global-model updates as momentum. At each round, the server keeps a momentum vector formed by accumulating the difference between consecutive global models. FedAdam [24] treats update differences as gradients optimized by Adam. Recently, regularization-based methods such as $\ell_{1}$-Fed [25], which alleviates similar distribution through the addition of an $\ell_{1}$ sparsity constraint to the global model, have also been reported. And FedGF [14], which resolves the difference between the local and global objectives by raising the generalization ability with Sharpness-Aware Minimization [6].
@@ 2 Related Works
When we construct a heterogeneous distribution with a simple semantics, such as a single target label, unified strategies exist to impose heterogeneity across clients by diversifying the prior distribution of the target label [31]. Specifically, two main strategies include (i) sampling the prior distribution of each client from Dirichlet distribution [1], and (ii) chunking per-class data samples into multiple shards, where a fixed number of shards is allocated to each client, resulting in heterogeneity between clients [2, 18].
2.1 Panoptic Scene Graph Generation (PSG)
Scene graphs are crucial for scene understanding in computer vision tasks, representing objects (nodes), denoted by bounding boxes or pixel-wise segmentation, and predicates (relationships, edges) in a graph structure [21, 12, 19]. Predicting the bounding boxes and relationships between bounding boxes constitutes scene graph generation. The PSG task has been proposed by [28, 30], delving deeper into scene graph generation using panoptic segmentation masks instead of bounding boxes. The difference between PSG and classic scene graph generation is that PSG uses panoptic segmentation [11] masks rather than bounding boxes. Moreover, the scene graph generation tasks face the long-tailed problem [5, 22]. Positional relationships among objects constitute the majority of the predicates, leading to a visual relationship complexity of $\mathcal{O}(N^{2}R)$ for $N$ objects and $R$ predicates [4]. This exacerbates the long-tailed problem in SGG datasets, prompting various approaches, such as utilizing self-attention to characterize complex interactions, thereby facilitating the understanding of object and relation semantics [15], and estimating confidence scores and weighting high-uncertainty cases more heavily during training [17]. Also, the self-supervised local pseudo-attribute is utilized to reinforce tail-class representations [9].
@@ 2.1 Panoptic Scene Graph Generation (PSG)
To perform challenging vision tasks such as PSG, having more data typically leads to training better models. However, in reality, the photos held by different clients are unlikely to be similar, and collecting such data on the server for training poses a threat to data privacy. Despite this, no attempts have been made to apply FL to the PSG task, underscoring the necessity of this research. Moreover, the data heterogeneity issue in FL closely resembles the well-known long-tailed problem in scene graph generation tasks and real-world data [22].
3 A Benchmark Process for FL with Multi-Semantic Datasets
@@
For a given multi-semantic dataset, each data sample contains multiple annotations, i.e., $(x,\mathcal{Y})\in\mathcal{D}$, where $x$ is an input, $\mathcal{Y}=\{y_{1},\ldots,y_{M}\}$ is a set of multi-semantic labels, $M$ is the possible number of labels for each data sample, and $\mathcal{D}$ represents the dataset. Here, we introduce our benchmark process to distribute the multi-semantic data samples to clients with controllable semantic heterogeneity. The key steps are twofold: (i) discovering data clusters with different semantics and (ii) data partitioning with controllable semantic heterogeneity across clients.
3.1 Discovering Data Clusters: $K$-means Clustering of Category Tensor
Figure 2: Category Tensor K-means Clustering Pipeline. (a) Construct a graph using the objects, subjects, and predicates of each image. (b) Map each label into the super-classes of fine-grained labels. (c) Convert categorized relations (subject, object, and predicate) into the category tensor. (d) Transform every input image into a category tensor and perform K-Means Clustering.
@@ 3.1 Discovering Data Clusters: $K$-means Clustering of Category Tensor
For a given multi-semantic $\mathcal{Y}$, we transform it to category tensor $\mathcal{F}$ by allocating each label $y_{i}$ into an orthogonal axis of the tensor, i.e., $\mathcal{F}(\mathcal{Y})\in\mathbb{R}^{N_{1}\times\cdots\times N_{L}}$, where there are $N_{1},\ldots,N_{L}$ possible categories for each respective label of $\mathcal{Y}$. We then apply $K$-means Clustering on the collection of $\mathcal{F(Y)}_{1}^{|\mathcal{D}|}$ of overall dataset:
@@
where $n$ is the number of clusters that can be determined depending on the dataset and $\mathcal{C}_{i}$ indicates the collection of samples assigned to $i$th cluster. With the obtained clustering, we can transform each data sample $(x,\mathcal{Y})$ into $(x,\mathcal{C}_{i})$ to impose the cluster label with semantic information, which is a one-hot label with $1$ for the assigned cluster. As a result of the clustering process, we can perform the label-based partition while fully utilizing the multi-semantic information of each data sample with its corresponding cluster label $\mathcal{C}$. We present the overall pipeline to 2.
3.2 Data Partition with Semantic Heterogeneity
We acquire $n$ clusters from Eq. (1). It trivially raises the issue that the clusters are not evenly distributed, so the number of samples assigned to each cluster would deviate for different clusters, i.e., cluster imbalance. The cluster imbalance prevents rigorous evaluations of FL models to handle semantic heterogeneity because a model becomes overfitted to dominant clusters without balanced training across different semantics. The cluster imbalance stems from the long-tailed problem, a key challenge in scene graph generation datasets. In other words, we have to create data heterogeneity for FL, which further complicates distinguishing it from the long-tailed problem. If the amount of data in each cluster is equalized, the long-tailed problem can be effectively alleviated. Furthermore, considering the FL scenario, this cluster imbalance will likely bias the update of the global model in the direction of users belonging to the dominant cluster. It causes overfitting to a dominant cluster, which makes it difficult to fairly compare each method closely. Consequently, we need to equalize the data quantity of each cluster: $\hat{\mathcal{C}}_{k}=\textnormal{Sample}(\mathcal{C}_{k},m)$, for all $1\leq k\leq n$, where $m=\min_{k\in[n]}\{|\mathcal{C}_{k}|\}$, $|\mathcal{C}_{k}|$ is the cardinality of the $k$th cluster $\mathcal{C}_{k}$, and $\textnormal{Sample}(\mathcal{C}_{k},m)$ functions to randomly select $m$ data samples from cluster $\mathcal{C}_{k}$. We apply the label-based partition based on these clusters to impose semantic heterogeneity. Our benchmark suggests two partition strategies as follows.